Suppose T "\isin" L(R2) is dened by T(x,y) = (-3y,x). Find the eigenvalues of T.


1
Expert's answer
2021-07-08T05:08:08-0400

Given that T is a linear transformation

T(x,y)=(-3y,x)

we know that the basis of R2 is {(1,0),(0,1)}

then,

T(1,0)=(-3(0),1)

T(1,0)=(0,1)

T(1,0)=0(1,0)+1(0,1)...........(1)

And,

T(0,1)=((-3)1,0)=(-3,0)

(0,1)=-3(1,0)+0(0,1)............(2)

Now the matrix representation of given linear transformation with respect to standard basis is given as,

[T]="\begin{bmatrix}
 0 & -3 \\
 1 & 0
\end{bmatrix}"

to find the eigen values of T we need to find the eigen values of matrix [T]

let A=[T]

then, A="\begin{bmatrix}
 0 & -3 \\
 1 & 0
\end{bmatrix}"

for eigen values put |A-"\lambda" I|=0

where"\lambda" is eigen value of A

then,

"\begin{bmatrix}
 0 & -3 \\
 1 & 0
\end{bmatrix}" -"\begin{bmatrix}
 \lambda & 0 \\
 0 & \lambda
\end{bmatrix}" =A-"\lambda" I

A-"\lambda" I="\begin{bmatrix}
 -\lambda & -3 \\
 1 & -\lambda
\end{bmatrix}"

put A-"\lambda" I=0

then,

"\begin{bmatrix}
 -\lambda & -3 \\
 1 & -\lambda
\end{bmatrix}" =0

"\lambda^2" +3=0

"\lambda^2" =-3

"\lambda=\sqrt3i,-\sqrt3i"

"\therefore" The eigen values of T are "\sqrt3i,-\sqrt3i"